OpenStudy (kaylardz0405):

i wil medal and fan solve 1/6 * 2 * 5 / 15

OpenStudy (anonymous):

\[Is ~this~your ~statement?\] \[\frac{ 1 }{ 6*2 }*\frac{ 5 }{15 }\]

OpenStudy (kaylardz0405):

the first fraction is right but the second is 5 divided by 15

OpenStudy (anonymous):

i also wrote 5/15

OpenStudy (kaylardz0405):

ok but 1/6 is a fraction

OpenStudy (anonymous):

1 divided by 6 or 1/6 is same

OpenStudy (kaylardz0405):

ok

OpenStudy (anonymous):

that means 2 is in the numerator.

OpenStudy (kaylardz0405):

opps was doing wrong question thanks anway

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }*2*\frac{ 5 }{ 15 }=\frac{ 2*5 }{ 6*15 }=\frac{ 1 }{ 3*3 }=?\]